Solve the system by substitution. x + y = 12
y = 2x

Answer by asha(30)

x+y =12 eqn,1
y =2x eqn.2
plugging the value of y in eqn.1 we get
x +2x =12
3x =12
x =12/3 =4
y =2x =2*4 =8
the answer is x=4 and y=8
good luck !!!